# Dimension-independent convergence rate of propagation of chaos and numerical analysis for McKean-Vlasov stochastic differential equations with coefficients nonlinearly dependent on measure

### v1

@@ Abstract
In contrast to ordinary stochastic differential equations (SDEs), the numerical simulation of McKean-Vlasov stochastic differential equations (MV-SDEs) requires approximating the distribution law first. Based on the theory of propagation of chaos, particle approximation method is widely used. Then, a natural question is to investigate the convergence rate of the method (also referred to as the convergence rate of PoC). In fact, the PoC convergence rate is well understood for MV-SDEs with coefficients linearly dependent on the measure, but the rate deteriorates with dimension $d$ under the $L^{p}$-Wasserstein metric for nonlinear measure-dependent coefficients, even when Lipschitz continuity with respect to the measure is assumed. The main objective of this paper is to establish a dimension-independent convergence result of PoC for MV-SDEs whose coefficients are nonlinear with respect to the measure component but Lipschitz continuous. As a complement we further give the time discretization of the equations and thus verify the convergence rate of PoC using numerical experiments.
Keywords:  McKean-Vlasov stochastic differential equations (MV-SDEs); Propagation of chaos (PoC); Tamed Euler-Maruyama(EM) method; Convergence rate; Dimension-independent.
@@ 1 Introduction
where $\mu_{t}^{x}$ denotes the law of the process $x$ at time $t$, i.e., $\mu_{t}^{x}=\mathbb{P}\circ x_{t}^{-1}$. To simulate this system, we expand it by introducing $N$ independent copies of $(x_{0},W)$, each generating a particle through
@@
${\rm d}X_{t}^{i,N}=f(X_{t}^{i,N},\mu_{t}^{X,N}){\rm d}t+g(X_{t}^{i,N},\mu_{t}^{X,N}){\rm d}W^{i}_{t},\quad X_{0}^{i,N}=x_{0}^{i},$ (3)
where $\mu_{t}^{X,N}=\frac{1}{N}\sum_{j=1}^{N}\delta_{X_{t}^{j,N}}$ is the empirical measure. As shown in [7], the empirical distribution $\mu_{t}^{X,N}$ converges to $\mu_{t}^{x^{i}}$ (and by uniqueness $\mu_{t}^{x^{i}}=\mu_{t}^{x}$), enabling approximation of the true measure. With this particle approximation in hand, discretizing (3) in time yields a numerical solution for the MV-SDEs (1).
@@ 2.1 Notations
Throughout this paper, let $(\Omega,\mathcal{F},\left\{\mathcal{F}_{t}\right\}_{t\geq 0},\mathbb{P})$ be a complete probability space with a filtration $\left\{\mathcal{F}_{t}\right\}_{t\geq 0}$ satisfying the usual conditions (i.e., it is right continuous and $\mathcal{F}_{0}$ contains all $\mathbb{P}$-null sets), and let $\mathbb{E}$ denote the expectation corresponding to $\mathbb{P}$. Let $|x|$ denote the Euclidean vector norm and $\langle x,y\rangle$ denote the inner product of vectors $x,y$. If $A$ is a vector or matrix, its transpose is denoted by $A^{\rm T}$. If A is a matrix, let $|A|$ denote its trace norm, i.e., $|A|=\sqrt{{\rm trace}(A^{\rm T}A)}$. In addition, we use $\mathcal{P}(\mathbb{R}^{d})$ to denote the family of all probability measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, where $\mathcal{B}(\mathbb{R}^{d})$ denotes the Borel $\sigma$-field over $\mathbb{R}^{d}$, and define the subset of probability measures with finite $p$-th moment by
@@
where $\mathcal{C}(\mu,\sigma)$ means all the couplings of $\mu$ and $\sigma$, i.e., $\pi\in\mathcal{C}(\mu,\sigma)$ if and only if $\pi(\cdot,\mathbb{R}^{d})=\mu(\cdot)$ and $\pi(\mathbb{R}^{d},\cdot)=\sigma(\cdot)$. For $t\in[0,T]$, $p>0$, $L_{t}^{p}(\mathbb{R}^{d})$ denotes the family of $\mathbb{R}^{d}$-valued $\mathcal{F}_{t}$-measurable random variables $\xi$ with $\mathbb{E}|\xi|^{p}<\infty$. As usual notations, “i.i.d.” means “independent and identically distributed”. In this paper we will often use the following fundamental inequality
$\left(\sum_{i=1}^{n}|a_{i}|\right)^{r}\leq n^{r-1}\sum_{i=1}^{n}|a_{i}|^{r},r\geq 1.$ (4)
Throughout this paper, $C$ denotes a generic positive constant that may change from line to line, and may depend on the initial data $x_{0}$ and the terminal time $T$ (and other fixed data), but is always independent of the dimension $d$ and the constants $\Delta$ and $N$ (associated with the numerical scheme and specified below).
@@ 2.2 Problem formulation
Let $x_{0}\in L_{0}^{p}(\mathbb{R}^{d})$ for some given $p\geq 2$ and independent of the $m$-dimensional Brownian motion $W$, $(x_{0}^{i},W^{i}),i=1,2,\dots,N$ are independent copies of $(x_{0},W)$. In the following, we work on the nonlinear MV-SDEs: for $i=1,2,\dots,N$,
${\rm d}x_{t}^{i}=$ $a(x_{t}^{i}){\rm d}t+f(x_{t}^{i},\mu_{t}^{x^{i}}){\rm d}t+g(x_{t}^{i},\mu_{t}^{x^{i}}){\rm d}W^{i}_{t},~{}t>0$ (5)
with
$f(x,\mu)=A\left(\int_{\mathbb{R}^{d}}\kappa(x,y)\mu({\rm d}y)\right),~{}g(x,\mu)=B\left(\int_{\mathbb{R}^{d}}\zeta(x,y)\mu({\rm d}y)\right)$ (6)
where $\mu_{t}^{x}$ denotes the law of the process $x$ at time $t$, i.e., $\mu_{t}^{x}=\mathbb{P}\circ x_{t}^{-1}$, and the interacting particle system
${\rm d}X_{t}^{i,N,\delta}=$ $\tilde{a}(X_{\delta_{t}}^{i,N,\delta}){\rm d}t+f(X_{\delta_{t}}^{i,N,\delta},\mu_{\delta_{t}}^{X,N,\delta}){\rm d}t+g(X_{\delta_{t}}^{i,N,\delta},\mu_{\delta_{t}}^{X,N,\delta}){\rm d}W^{i}_{t},~{}t>0$ (7)
@@
where $\delta_{t}=[t/\delta]\delta$ for $\delta>0$, $[t/\delta]$ denotes the integer part of $t/\delta$. $\mu_{t}^{X,N,\delta}=\frac{1}{N}\sum_{j=1}^{N}\delta_{X_{t}^{j,N,\delta}}$, $a,\tilde{a}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $A:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, $B:\mathbb{R}^{d\times m}\to\mathbb{R}^{d\times m}$, $\zeta:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$.
The main aim in this section is to establish the following estimate:
$\mathbb{E}|x_{t}^{i}-X_{t}^{i,N,\delta}|^{p}\leq$ $C\left\{N^{-p/2}+(t-\delta_{t})^{p/2}+\mathbb{E}|a(X_{t}^{i,N,\delta})-\tilde{a}(X_{t}^{i,N,\delta})|^{p}\right\}.$ (8)
@@
for $t\in[0,T]$ with $T$ is the given terminal time. Next, let us explain why we focus on the systems (5) and (7), and consider the estimate (8). This is primarily based on the following facts:
@@
    ${\rm d}X_{t}^{i,N}=$ $a(X_{t}^{i,N}){\rm d}t+f(X_{t}^{i,N},\mu_{t}^{X,N}){\rm d}t+g(X_{t}^{i,N},\mu_{t}^{X,N}){\rm d}W^{i}_{t},$ (9)
    where we set $X_{t}^{i,N}:=X_{t}^{i,N,t},\mu_{t}^{X,N}:=\mu_{t}^{X,N,t}$ for the notational brevity. It is easy to see that (5) and (9) are the non-interacting particle system and the interacting particle system corresponding to the MV-SDEs
@@
    respectively, and as an immediate by-product of (8), the PoC result
    $\mathbb{W}_{p}(\mu_{t}^{x^{i}},\mu_{t}^{X^{i,N}})\leq\left(\mathbb{E}|x_{t}^{i}-X_{t}^{i,N}|^{p}\right)^{1/p}\leq$ $CN^{-1/2}$
@@
  • If the coefficients of equation (5) are all linearly growing, it is well known that the EM scheme is the simplest method to simulate the MV-SDEs: for a step size $\Delta>0$, $\underline{t}:=[t/\Delta]\Delta$,
    ${\rm d}X_{t}^{i,N,\Delta}=$ $a(X_{\underline{t}}^{i,N,\Delta}){\rm d}t+f(X_{\underline{t}}^{i,N,\Delta},\mu_{\underline{t}}^{X,N,\Delta}){\rm d}t+g(X_{\underline{t}}^{i,N,\Delta},\mu_{\underline{t}}^{X,N,\Delta}){\rm d}W^{i}_{t}.$
@@
    where the precise definitions of the functions $\varphi$ and $h$ can be found in [18], then (7) becomes the tamed EM and truncated EM schemes for system (5), respectively, and (8) gives the convergence result of the numerical scheme, as long as we further give the error estimate between $a$ and $\tilde{a}$.
@@
2.3 Error estimate between (5) and (7)
Before starting our work, the first natural thing to do is to ensure that equations (5) and (7) have unique solutions, we do these under the following settings.
@@ 2.3 Error estimate between (5) and (7)
There is a constant $L>0$ such that
$|a(0)|\vee|A(0)|\vee|\kappa(0,0)|\vee|B(0)|\vee|\zeta(0,0)|\leq L.$
Assumption 2.2.
There are constants $\lambda_{1}\in\mathbb{R}$ and $L_{1},q>0$ such that
$\langle x-y,a(x)-a(y)\rangle\leq$ $\lambda_{1}|x-y|^{2},$
$|a(x)-a(y)|\leq$ $L_{1}(1+|x|^{q}+|y|^{q})|x-y|$
@@
There are constants $\lambda_{2},\lambda_{3},L_{2},L_{3}>0$ such that
$|A(x)-A(y)|\leq$ $\lambda_{2}|x-y|,\quad|\kappa(x,y)-\kappa(\bar{x},\bar{y})|\leq L_{2}(|x-\bar{x}|+|y-\bar{y}|),$
$|B(x)-B(y)|\leq$ $\lambda_{3}|x-y|,\quad|\zeta(x,y)-\zeta(\bar{x},\bar{y})|\leq L_{3}(|x-\bar{x}|+|y-\bar{y}|)$
@@
It is well known that Assumption 2.1 imply the following growth conditions of the coefficients:
$\langle x,a(x)\rangle=$ $\langle x-0,a(x)-a(0)\rangle+\langle x,a(0)\rangle\leq\lambda_{1}|x|^{2}+\frac{|x|^{2}}{2}+\frac{|a(0)|^{2}}{2}\leq(\lambda_{1}+\frac{1}{2})|x|^{2}+\frac{L^{2}}{2},$
$|a(x)|\leq$ $|a(x)-a(0)|+|a(0)|\leq L_{1}(|x|+|x|^{q+1})+L,$
@@
$|A(x)|\leq$ $\lambda_{2}|x|+L,~{}|\kappa(x,y)|\leq L_{2}(|x|+|y|)+L,$
$|B(x)|\leq$ $\lambda_{3}|x|+L,~{}|\zeta(x,y)|\leq L_{3}(|x|+|y|)+L$
@@
Let $p\geq 2$, for any $i=1,2,\dots,N$, under Assumption 2.1, there exists a unique solution $x_{t}^{i}$ to the equation (5), and the solution satisfies

### v2

@@ Abstract
In contrast to ordinary stochastic differential equations (SDEs), the numerical simulation of McKean-Vlasov stochastic differential equations (MV-SDEs) requires approximating the distribution law first. Based on the theory of propagation of chaos (PoC), particle approximation method is widely used. Then, a natural question is to investigate the convergence rate of the method (also referred to as the convergence rate of PoC). In fact, the PoC convergence rate is well understood for MV-SDEs with coefficients linearly dependent on the measure, but the rate deteriorates with dimension $d$ under the $L^{p}$-Wasserstein metric for nonlinear measure-dependent coefficients, even when Lipschitz continuity with respect to the measure is assumed. The main objective of this paper is to establish a dimension-independent convergence result of PoC for MV-SDEs whose coefficients are nonlinear with respect to the measure component but Lipschitz continuous. As a complement we further give the time discretization of the equations and thus verify the convergence rate of PoC using numerical experiments.
Keywords:  McKean-Vlasov stochastic differential equations (MV-SDEs); Propagation of chaos (PoC); Convergence rate; Dimension-independent; Tamed Euler-Maruyama (EM) method.
@@ 1 Introduction
with the initial data $x_{0}$, where $\mu_{t}^{x}$ denotes the law of the process $x$ at time $t$, i.e., $\mu_{t}^{x}=\mathbb{P}\circ x_{t}^{-1}$. To simulate this system, let $(x_{0}^{i},W^{i})_{1\leq i\leq N}$ be $N$ independent copies of $(x_{0},W)$, each generating a particle through
@@
${\rm d}X_{t}^{i,N}=f(X_{t}^{i,N},\hat{\mu}_{t}^{X,N}){\rm d}t+g(X_{t}^{i,N},\hat{\mu}_{t}^{X,N}){\rm d}W^{i}_{t},\quad X_{0}^{i,N}=x_{0}^{i},$ (3)
where $\hat{\mu}_{t}^{X,N}=\frac{1}{N}\sum_{j=1}^{N}\delta_{X_{t}^{j,N}}$ is the empirical measure. As shown in [7], the empirical distribution $\hat{\mu}_{t}^{X,N}$ converges to $\mu_{t}^{x^{i}}$ (and by uniqueness $\mu_{t}^{x^{i}}=\mu_{t}^{x}$), enabling approximation of the true measure. With this particle approximation in hand, discretizing (3) in time yields a numerical solution for the MV-SDEs (1).
@@ 2.1 Notations
Throughout this paper, let $(\Omega,\mathcal{F},\left\{\mathcal{F}_{t}\right\}_{t\geq 0},\mathbb{P})$ be a complete probability space with a filtration $\left\{\mathcal{F}_{t}\right\}_{t\geq 0}$ satisfying the usual conditions (i.e., it is right continuous and $\mathcal{F}_{0}$ contains all $\mathbb{P}$-null sets), and let $\mathbb{E}$ denote the expectation corresponding to $\mathbb{P}$. Let $|x|$ denote the Euclidean vector norm and $\langle x,y\rangle$ denote the inner product of vectors $x,y$. If $A$ is a vector or matrix, its transpose is denoted by $A^{\rm T}$. If A is a matrix, let $|A|$ denote its trace norm, i.e., $|A|=\sqrt{{\rm trace}(A^{\rm T}A)}$. For $a,b\in\mathbb{R},a\vee b:=\max\{a,b\},a\wedge b:=\min\{a,b\}$. In addition, we use $\mathcal{P}(\mathbb{R}^{d})$ to denote the family of all probability measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, where $\mathcal{B}(\mathbb{R}^{d})$ denotes the Borel $\sigma$-field over $\mathbb{R}^{d}$, and define the subset of probability measures with finite $p$-th moment by
@@
where $\mathcal{C}(\mu,\sigma)$ means all the couplings of $\mu$ and $\sigma$, i.e., $\pi\in\mathcal{C}(\mu,\sigma)$ if and only if $\pi(\cdot,\mathbb{R}^{d})=\mu(\cdot)$ and $\pi(\mathbb{R}^{d},\cdot)=\sigma(\cdot)$. For $t\in[0,T]$, $p>0$, $L_{t}^{p}(\mathbb{R}^{d})$ denotes the family of $\mathbb{R}^{d}$-valued $\mathcal{F}_{t}$-measurable random variables $\xi$ with $\mathbb{E}|\xi|^{p}<\infty$. As usual notations, “i.i.d.” means “independent and identically distributed”.
@@ 2.2 Problem formulation
Let $x_{0}\in L_{0}^{p}(\mathbb{R}^{d})$ for some given $p\geq 2$ and independent of the $m_{0}$-dimensional Brownian motion $W$. Let $T>0$, $a:\mathbb{R}^{d}\to\mathbb{R}^{d}$, in the following, we work on the nonlinear MV-SDEs: for $i=1,2,\dots,N$,
${\rm d}x_{t}^{i}=$ $a(x_{t}^{i}){\rm d}t+f(x_{t}^{i},\mu_{t}^{x^{i}}){\rm d}t+g(x_{t}^{i},\mu_{t}^{x^{i}}){\rm d}W^{i}_{t},~{}t\in[0,T],$ (4)
where $(x_{0}^{i},W^{i})_{1\leq i\leq N}$ are independent copies of $(x_{0},W)$. For $t\in[0,T]$, define $\delta_{t}=[t/\delta]\delta$ for $\delta>0$ with $[t/\delta]$ denotes the integer part of $t/\delta$, $\hat{\mu}_{t}^{X,N,\delta}=\frac{1}{N}\sum_{j=1}^{N}\delta_{X_{t}^{j,N,\delta}}$. Let $X_{0}^{i,N,\delta}=x_{0}^{i}$ for $i=1,\dots,N$, $\tilde{a}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, we also consider the interacting particle system
${\rm d}X_{t}^{i,N,\delta}=$ $\tilde{a}(X_{\delta_{t}}^{i,N,\delta}){\rm d}t+f(X_{\delta_{t}}^{i,N,\delta},\hat{\mu}_{\delta_{t}}^{X,N,\delta}){\rm d}t+g(X_{\delta_{t}}^{i,N,\delta},\hat{\mu}_{\delta_{t}}^{X,N,\delta}){\rm d}W^{i}_{t},~{}t\in[0,T]$ (5)
with coefficients taking the form of
  • (i)
    nonlinear distribution-dependence case:
    $f(x,\mu)=A\left(\int_{\mathbb{R}^{d}}\kappa(x,y)\mu({\rm d}y)\right),~{}g(x,\mu)=B\left(\int_{\mathbb{R}^{d}}\zeta(x,y)\mu({\rm d}y)\right),$ (6)
    where $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, $A:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $\zeta:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m_{0}}$, $B:\mathbb{R}^{d\times m_{0}}\to\mathbb{R}^{d\times m_{0}}$.
  • (ii)
    more general nonlinear dependence case:
    $f(x,\mu)=$ $A\left(\int_{\mathbb{R}^{d}}\kappa_{1}(x,y)\mu({\rm d}y),\cdots,\int_{\mathbb{R}^{d}}\kappa_{q}(x,y)\mu({\rm d}y)\right),$ (7)
    $g(x,\mu)=$ $B\left(\int_{\mathbb{R}^{d}}\zeta_{1}(x,y)\mu({\rm d}y),\cdots,\int_{\mathbb{R}^{d}}\zeta_{q}(x,y)\mu({\rm d}y)\right)$ (8)
@@
    where $q$ is a positive integer, $\kappa_{j}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d},\zeta_{j}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m_{0}},j=1,\dots,q$, $A:\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, $B:\mathbb{R}^{d\times m_{0}}\times\cdots\times\mathbb{R}^{d\times m_{0}}\to\mathbb{R}^{d\times m_{0}}$.
  • (iii)
    higher-order interaction case:
    $f(x,\mu)=$ $A\left(\int_{\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}}\kappa(x,y_{1},\dots,y_{q})\mu({\rm d}y_{1})\cdots\mu({\rm d}y_{q})\right),$ (9)
    $g(x,\mu)=$ $B\left(\int_{\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}}\zeta(x,y_{1},\dots,y_{q})\mu({\rm d}y_{1})\cdots\mu({\rm d}y_{q})\right),$ (10)
    where $q$ is a positive integer, $\kappa:\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, $\zeta:\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m_{0}},$ $A:\mathbb{R}^{d}\to\mathbb{R}^{d},B:\mathbb{R}^{d\times m_{0}}\to\mathbb{R}^{d\times m_{0}}$.
Below, we make a comment on the three different forms of the coefficients.
Remark 2.1.
Note that (6) does not include the cases when several statistical quantities are involved in the coefficients, for example, if the coefficient depends on the variance, then both the first and second moments will be involved. These cases could be given by (7) and (8) by introducing multiple kernel functions. On the other hand, if the system depends on higher-order statistics, such as covariance, it is necessary to be described using equations with higher-order interactions (9) and (10).
The main aim in section 3 is to establish the following error estimate:
$\max_{1\leq i\leq N}\mathbb{E}|x_{t}^{i}-X_{t}^{i,N,\delta}|^{p}\leq$ $C\left\{N^{-\frac{p}{2}}+(t-\delta_{t})^{\frac{p}{2}}+\max_{1\leq i\leq N}\mathbb{E}|a(X_{t}^{i,N,\delta})-\tilde{a}(X_{t}^{i,N,\delta})|^{p}\right\}$ (11)
for $p\geq 2$. Throughout this paper, let $C$ denote a generic positive constant that may change from line to line, and may depend on the initial data $x_{0}$, the terminal time $T$, and other fixed data (such as $p$ and the Lipschitz constants introduced below), but is always independent of the dimension $d$, the number of particles $N$ and the constant $\delta$.
@@
Next, let us explain why we focus on the systems (4) and (5), and consider the estimate (11). This is primarily based on the following facts:
@@
    ${\rm d}X_{t}^{i,N}=$ $a(X_{t}^{i,N}){\rm d}t+f(X_{t}^{i,N},\hat{\mu}_{t}^{X,N}){\rm d}t+g(X_{t}^{i,N},\hat{\mu}_{t}^{X,N}){\rm d}W^{i}_{t},~{}X_{0}^{i,N}=x_{0}^{i},$ (12)
    where we set $X_{t}^{i,N}:=X_{t}^{i,N,t},\hat{\mu}_{t}^{X,N}:=\hat{\mu}_{t}^{X,N,t}$ for the notational brevity. It is easy to see that (4) and (12) are the non-interacting particle system and the interacting particle system corresponding to the MV-SDEs
@@
    with the initial data $x_{0}$, respectively, and as an immediate by-product of (11), the PoC result
    $\max_{1\leq i\leq N}\left\{\left(\mathbb{E}|x_{t}^{i}-X_{t}^{i,N}|^{p}\right)^{1/p}+\mathbb{E}\mathbb{W}_{p}(\mu_{t}^{x^{i}},\hat{\mu}_{t}^{X,N})\right\}\leq CN^{-1/2}$
@@
  • If the coefficients of equation (4) are all linearly growing, it is well known that the EM scheme is the simplest method to simulate the MV-SDEs ([26]): for a step size $\Delta>0$, $\underline{t}:=[t/\Delta]\Delta$,
    ${\rm d}X_{t}^{i,N,\Delta}=$ $a(X_{\underline{t}}^{i,N,\Delta}){\rm d}t+f(X_{\underline{t}}^{i,N,\Delta},\hat{\mu}_{\underline{t}}^{X,N,\Delta}){\rm d}t+g(X_{\underline{t}}^{i,N,\Delta},\hat{\mu}_{\underline{t}}^{X,N,\Delta}){\rm d}W^{i}_{t}.$
@@
    where the precise definitions of the functions $\varphi$ and $h$ can be found in [18, Section 3], then (5) becomes the tamed EM and truncated EM schemes for system (4), respectively, and (11) gives the convergence result of the numerical scheme, as long as we further give the error estimate between $a$ and $\tilde{a}$.
@@
3 Main results
3.1 Error estimate under the nonlinear distribution-dependence case
The main objective of this paper is to establish the optimal convergence rate of the numerical method, particularly for the propagation of chaos, when the coefficients nonlinearly depend on the measure components. First, we consider the equation (4) with coefficients taking the form of (6). Before starting our work, the first natural step is to ensure that equations (4) and (5) have unique solutions in this setting. We establish this under the following conditions.
@@ 3.1 Error estimate under the nonlinear distribution-dependence case
There are constants $L_{1}>0$ and $r\geq 1$ such that
$\langle x-y,a(x)-a(y)\rangle\leq$ $L_{1}|x-y|^{2},$
$|a(x)-a(y)|\leq$ $L_{1}(1+|x|^{r}+|y|^{r})|x-y|,$
$|a(0)|\leq$ $L_{1}$
@@
There are constants $L_{2}>0$ such that
$|A(x)-A(y)|\vee|B(x)-B(y)|\leq$ $L_{2}|x-y|,$
$|\kappa(x,y)-\kappa(\bar{x},\bar{y})|\vee|\zeta(x,y)-\zeta(\bar{x},\bar{y})|\leq$ $L_{2}(|x-\bar{x}|+|y-\bar{y}|),$
$|A(0)|\vee|\kappa(0,0)|\vee|B(0)|\vee|\zeta(0,0)|\leq$ $L_{2}$
@@
$\sup_{1\leq i\leq N}\left\{\mathbb{E}\left(\sup_{0\leq t\leq T}|X_{t}^{i,N,\delta}|^{p}\right)\vee\mathbb{E}\left(\sup_{0\leq t\leq T}|\tilde{a}(X_{t}^{i,N,\delta})|^{p}\right)\right\}\leq$ $C,$
where $C$ is independent of $d$, $N$ and $\delta$.
@@
It is well known that Assumptions 3.1 and 3.2 imply the following growth conditions of the coefficients:
$\langle x,a(x)\rangle=$ $\langle x-0,a(x)-a(0)\rangle+\langle x,a(0)\rangle\leq L_{1}|x|^{2}+\frac{|x|^{2}}{2}+\frac{|a(0)|^{2}}{2}\leq(L_{1}+\frac{1}{2})|x|^{2}+\frac{L_{1}^{2}}{2},$
$|a(x)|\leq$ $|a(x)-a(0)|+|a(0)|\leq L_{1}(1+|x|+|x|^{r+1}),$
@@
$|A(x)|\vee|B(x)|\leq$ $L_{2}(1+|x|),~{}|\zeta(x,y)|\vee|\kappa(x,y)|\leq L_{2}(1+|x|+|y|),$
@@
Let $p\geq 2$, for any $i=1,2,\dots,N$, under Assumptions 3.1 and 3.2, there exists a unique solution $x_{t}^{i}$ to the equation (4), and the solution satisfies